Write an equation for the line described. Write the equation in the form specified.


parallel to y + 4x = 4, through (4, 4); slope-intercept form

Answer by ikleyn(52781)

.
Write an equation for the line described. Write the equation in the form specified.
parallel to y + 4x = 4, through (4, 4); slope-intercept form
~~~~~~~~~~~~~~~~~~~~~~~~~~~

Line parallel to y + 4x = 4  has an equation y + 4x = c,

where "c" is some constant.  To find it, substitute the coordinates of the point (4,4)
into this equation.  You will get  c = 4 + 4*4 = 20.


Thus, the line has an equation y + 4x = 20.


In the slope-intercept form, this equation is

    y = -4x + 20.    ANSWER

Solved.